# SWIFT:A Scalable lightWeight Infrastructure for Fine-Tuning (work in progress)

### v3

@@ 3.1.2. Evaluation Architecture Design
For evaluation capabilities, we rely on the EvalScope171717GitHub:https://github.com/modelscope/evalscope framework from the ModelScope Community. This framework constructs evaluation capabilities by integrating OpenCompass (for text models) and VLMEvalKit(duan2024vlmevalkit, ) (for multi-modal models). By incorporating EvalScope, SWIFT currently supports over 100+ total pure text evaluation sets and multi-modal evaluation sets, as well as the aforementioned two types of custom evaluation datasets and their evaluation processes.
@@ 3. API Calling
Based on LoRA training, the average metrics of LLaMA3 improved by 17%. This demonstrates that open-source models and datasets are meaningful for Agent training in practical vertical scenarios. We have summarized the hyper-parameters and other experiences from the training process to facilitate replication and application by other developers. The mentioned dataset 181818https://www.modelscope.cn/datasets/iic/MSAgent-Pro and models 191919WebPage: https://modelscope.cn/models/swift/qwen2-7b-agent-instruct 202020WebPage: https://modelscope.cn/models/swift/llama3-8b-agent-instruct-v2 can all be found on ModelScope.

### v4

@@ 3.1.2. Evaluation Architecture Design
For evaluation capabilities, we rely on the EvalScope191919GitHub:https://github.com/modelscope/evalscope framework from the ModelScope Community. This framework constructs evaluation capabilities by integrating OpenCompass (for text models) and VLMEvalKit(duan2024vlmevalkit, ) (for multi-modal models). By incorporating EvalScope, SWIFT currently supports over 100+ total pure text evaluation sets and multi-modal evaluation sets, as well as the aforementioned two types of custom evaluation datasets and their evaluation processes.
@@ 3. API Calling
Based on LoRA training, the average metrics of LLaMA3 improved by 17%. This demonstrates that open-source models and datasets are meaningful for Agent training in practical vertical scenarios. We have summarized the hyper-parameters and other experiences from the training process to facilitate replication and application by other developers. The mentioned dataset 202020https://www.modelscope.cn/datasets/iic/MSAgent-Pro and models 212121WebPage: https://modelscope.cn/models/swift/qwen2-7b-agent-instruct 222222WebPage: https://modelscope.cn/models/swift/llama3-8b-agent-instruct-v2 can all be found on ModelScope.